OpenStudy (anonymous):

what is the gcf of -30a^3, 5a^4

OpenStudy (anonymous):

\[5a^3\]

OpenStudy (anonymous):

since it is both a factor of \[5a^4\] via \[5a^3\times a = 5a^4\] and also a factor of \[-30x^3\] since \[-6\times 5a^3=-30a^3\]